. A 100ml water sample on titration with N/50H2SO4required 4.7 ml of acid upto phenolphthalein endpoint. Addition of methyl orange to the same solution required 10.5ml of same acid. Elucidate the type and extent of alkalinity present in water sample.


1
Expert's answer
2021-06-29T00:39:01-0400

Here diacidic base is present which give two end point one with phenalphthein and second with methy orange

For phenolphthalein end point-

"\frac{1}{2}\ Meq\ of\ Diacidic\ base = Meq.\ of\ of \ 8ml \ \frac{N}{50}\ HCl"

="\frac{1\times 8}{50\times1000}=\frac{1}{6225}\ Meq\\"

"Meq.\ of\ Diacidic\ base=2\times Meq.\ of\ HCl=\frac{2}{6225}\ Meq."

For methyl orange end point-

"Meq.\ of\ monoacidic\ base\ formed=Meq.\ of\ 1ml of\ extra\ HCl\ used"

"=\frac{1\times1}{50\times 1000}=\frac{1}{50000\ Meq.}"


Normality of diacidic alkalinity present in 100 ml water = "\frac{2\times 1000}{6225\times 100}=\frac{20}{6225}=\frac{4}{1245}\ N"